Compute for the concentration of perchloric acid solution if the pOH = 5.52


1
Expert's answer
2021-02-15T03:26:04-0500

pH = 14 - pOH = 14 - 5.52 = 8.48

There is a mistake in the task: pH of the strong acid can't be more than 7

So, if there should be pH = 5.52 in the task, the the solution:

pH = -lg(H+)

[H+] = 10-pH = 10-5.52 = 3.02 × 10-6

HClO4 is a strong acid, which means CHClO4 = [H+] = 3.02 × 10-6 (M)

Answer: CHClO4 = 3.02 × 10-6 M.